Impulse on a Baseball 12 of 15 > Constants The impulse of an object is also related to its change in momentum. Once the impulLearning Goal: To understand the relationship between force, impulse, and momentum. The effect of a net force ΣF⃗ acting on an object is related both to the force and to the total time the force acts on the object. The physical quantity impulse J⃗ is a measure of both these effects. For a constant net force, the impulse is given by J⃗ =F⃗ Δt. The impulse is a vector pointing in the same direction as the force vector. The units of J⃗ are N⋅s or kg⋅m/s. Recall that when a net force acts on an object, the object will accelerate, causing a change in its velocity. Hence the object's momentum (p⃗ =mv⃗ ) will also change. The impulse-momentum theorem describes the effect that an impulse has on an object's motion: Δp⃗ =J⃗ =F⃗ Δt. So the change in momentum of an object equals the net impulse, that is, the net force multiplied by the time over which the force acts. A given change in momentum can result from a large force over a short time or a smaller force over a longer time.

Please answer part e
Answer #1

J= Impulse

Vf=Final velocity

Vi=Initial velocity

m=Mass of the object

J=m(Vf-Vi/) =

Vf= J/m + Vi= (-8.4/0.145) + 32

Vf= -25.93 m/s

Final velocity of the ball is 25.93m/s in -x direction